OpenStudy (anonymous):

The length of a rectangle is 5 meters more than width. If the area is 84 square meters, what are the length and width?

OpenStudy (anonymous):

x(x+5) =84 ==> x^2+5x-84=0 ===> (x+12)(x-7)=0 ===> x=7 width ==> length =12